OpenStudy (anonymous):

Find the solution to the following equation by transforming it into a perfect square trinomial. x2 – 10x = 39 A.{–3, 13} B.{–15, 25} C.{–20, 30} D.{59, 69}

OpenStudy (anonymous):

x2 – 10x = 39 split -10 into two for -5 and -5...multiply them to get 25 and add 25 to both sides x2 – 10x +25= 39+25

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

x2 – 10x+25 = 64 then factor the left side and take the square root of both sides...

OpenStudy (anonymous):

hope that was enough help

OpenStudy (anonymous):

?

OpenStudy (anonymous):

was it? enough help?

OpenStudy (anonymous):

No... I don't get your second part

OpenStudy (anonymous):

can you factor the left side?

OpenStudy (anonymous):

Sure, what number should I factor?

OpenStudy (anonymous):

you need to factor the polynomial... Unfoil it...

OpenStudy (anonymous):

How do you factor a polynomial?

OpenStudy (anonymous):

\[(x-5)(x-5)\] do you know how to multiply those two...

OpenStudy (anonymous):

FOIL?

OpenStudy (anonymous):

x^2-5x-5x+25

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

yes...now combine like terms

OpenStudy (anonymous):

-5x-5x = -10x

OpenStudy (anonymous):

x^2-10x+25

OpenStudy (anonymous):

so \[(x-5)^2=64\]

OpenStudy (anonymous):

because \[x^2-10x+25=64=(x-5)(x-5)\]

OpenStudy (anonymous):

Ok...

OpenStudy (anonymous):

so take the square root of both sides of the equation \[(x-5)^2=64\]

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

and you get \[x-5=8\] or x-5=-8

OpenStudy (anonymous):

So...x=-3

OpenStudy (anonymous):

\[\sqrt{(x-5)^2}=(x-5)\] and \[\sqrt{64}=\pm8\]

OpenStudy (anonymous):

yes...x=-3 is one possible answer

OpenStudy (anonymous):

Isn't there a more simply way of solving this type of question?

OpenStudy (anonymous):

So the other x=13?

OpenStudy (anonymous):

yes...completing the square is one way to do it...

OpenStudy (anonymous):

the easier way is to simply factor the initial polynomial

OpenStudy (anonymous):

either way...you will need to learn how to factor polynomials...

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

multiply these two and combine like terms \[(x+3)(x-13)\]

OpenStudy (anonymous):

Ah, soo.. how would you factor the initial polynomial

OpenStudy (anonymous):

Using the FOIL method?

OpenStudy (anonymous):

Sorry, my OS is loading extremely slow

OpenStudy (anonymous):

\[x^2-10x=39\]

OpenStudy (anonymous):

mine is going slow too

OpenStudy (anonymous):

subtract 39 from both sides

OpenStudy (anonymous):

\[x^2-10x-39=0\]

OpenStudy (anonymous):

\[x^2-x13+3x-39\]

OpenStudy (anonymous):

factors to\[(x+3)(x-13)=0\] now either \[x+3=0\] or x-13=0

OpenStudy (anonymous):

https://www.youtube.com/watch?v=bTe-HhNkR28

OpenStudy (anonymous):

watch that:)

OpenStudy (anonymous):

ooo thanks

OpenStudy (anonymous):

yup:)